CH/DOC/94/07

Grant Funds

Proposed remaining spend

Author: B.R. Fulton Date: 21 April 1994

Proposed spend for remaining grant funds

Attached is a proposed spend for the residue of the money still in each groups grant. Each group will have to get orders placed in time to meet the deadline imposed by their University accounts office. To discuss and agree at meeting.

Item  Price (£)

Birmingham (Remaining funds £21,000)

Shipment to ANU in July  1,000
Refurbishment of NaI array  1,000
Replenish ANU account  5,000
Strip detectors for July  10,000
Shipment of Strasbourg beamline  5,000

Oxford (Remaining funds £80,000)

Final layout and components for amps  30,300
Production of amp crates  2,800
Components for walk/accept  9,000
Build walk/accept  2,200
Components for convertors  10,700
Components for 2 FETCH modules  4,100
Software for Starburst  3,500
Components for special modules  5,900
Advance spend on computing  11,000

(continued)

Item  Price (£)

Surrey (Remaining funds £61,000)

3 off 32 channel TDC’s  10,500
320 Silena sample and hold circuits  23,700
8 off Silena ADC  20,000
Camac Crate (for above)  5,000
LeCroy data stack (for above)  2,000

York (Remaining funds £70,000)

500 preamplifiers  6,000
80 bias supplies  1,000
MEGHA chamber  10,000
Connectors for chamber  4,000
Intertechnique  5,000
Ethernet hardware  8,000

(York figures given by DLW at Charissa Meeting of 27 April 1994)

Brian Fulton 21.4.94